OpenStudy (anonymous):

How far could you speed walk in 10 minutes based on your speed for the 10 meter trial? Time: 6.23 seconds Speed: 1.6 meter per seconds

OpenStudy (anonymous):

Speed = distance / time Speed = 1.6 m/s distance = speed * time 10 minutes = 10*60 s = 600 s distance = (1.6 m/s)*(600 s) distance = 9.6*10^2 m